OpenStudy (anonymous):

There's a function table and they want me to find the missing value. It looks like this: x y 0 -1 1 1 3 4 7 5 9 I need help figuring out what the function is between them also, if you have a method please share? Thanks :)

OpenStudy (johnweldon1993):

well lets look x goes from 0-1...then 3....then 4,5....so it looks like it would go up by 1 every time right? so lets write that x 0 1 2 3 4 5 okay now y start -1...goes to 1....what is that? a +2 change right? from 4-5....we go from 7 to 9 right? ALSO a +2 change... so we can assume it will be consistant and go up by 2 every number so -1 1 3 5 7 9 hmm it works out doesnt it.... so we have x y 0 -1 1 1 2 3 3 5 4 7 5 9 look right?

OpenStudy (anonymous):

Wait but what exactly would be the function? Shouldn't there be a set function?

OpenStudy (johnweldon1993):

yes....the function would be y = 2x - 1 makes sense right? if x is 0 2(0) - 1 = -1 x is 1 2(1) - 1 = 1 x=2 2(2) - 1 = 3 make sense?

OpenStudy (anonymous):

Yes, thank you so much for your replies.

OpenStudy (johnweldon1993):

not a problem :)